the diameter of a sphere is 6 cm. it is melted and drawn into a wire of diameter 0.2cm . find the length of the wire.

Answer by mananth(16946)

diameter of sphere = 6cm
radius =3
Vol of sphere = 4/3 * pi* r^3
Volume = 4/3 *pi*27
Volume = 113.09 cm^3
--
Volume of cylinder = pi*r^2*h
113.09=pi*(0.2)^2*l
l=113.09/(pi*(0.2)^2)
l= 899.94=> 900 cm
m.ananth@hotmail.ca